Determine graphically the x-value when y= -5 in the given equation
-x+2y= -8
The x-value when y= -5 is x ??

Answer by jim_thompson5910(35256)

-x+2y= -8


-x+2(-5)= -8 ... replace y with -5


-x-10 = -8 ... multiply


Now solve for x (I'll let you do this). Let me know if you still need help. Thanks.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------